# ConspEmoLLM: Conspiracy Theory Detection Using an Emotion-Based Large Language Model

### v1

@@ Abstract
The internet has brought both benefits and harms to society. A prime example of the latter is misinformation, including conspiracy theories, which flood the web. Recent advances in natural language processing, particularly the emergence of large language models (LLMs), have improved the prospects of accurate misinformation detection. However, most LLM-based approaches to conspiracy theory detection focus only on binary classification and fail to account for the important relationship between misinformation and affective features (i.e., sentiment and emotions). Driven by a comprehensive analysis of conspiracy text that reveals its distinctive affective features, we propose ConspEmoLLM, the first open-source LLM that integrates affective information and is able to perform diverse tasks relating to conspiracy theories. These tasks include not only conspiracy theory detection, but also classification of theory type and detection of related discussion (e.g., opinions towards theories). ConspEmoLLM is fine-tuned based on an emotion-oriented LLM using our novel ConDID dataset, which includes five tasks to support LLM instruction tuning and evaluation. We demonstrate that when applied to these tasks, ConspEmoLLM largely outperforms several open-source general domain LLMs and ChatGPT, as well as an LLM that has been fine-tuned using ConDID, but which does not use affective features. This project will be released at https://github.com/lzw108/ConspEmoLLM/.
ConspEmoLLM: Conspiracy Theory Detection Using an Emotion-Based Large Language Model
Zhiwei Liu, Boyang Liu, Paul Thompson, Kailai Yang, Raghav Jain, Sophia Ananiadou National Centre for Text Mining, Department of Computer Science The University of Manchester, Manchester, United Kingdom {zhiwei.liu-2, boyang.liu-2, kailai.yang}@postgrad.manchester.ac.uk {paul.thompson, raghav.jain, sophia.ananiadou}@manchester.ac.uk
@@ 1 Introduction
Misinformation has become one of the major threats to society. The rise of the internet and social media has made it increasingly simple for misinformation to spread rapidly. Conspiracy theories are one type of misinformation, whose false content is intended to cause harm Napolitano and Reuter (2023). Examples of popular conspiracy theories include those claiming that the Earth is flat and that vaccines cause autism. Conspiracy theorists ignore scientific evidence and tend to interpret events as secretive actions Giachanou et al. (2023). During the COVID-19 pandemic, the spread of conspiracy theories significantly increased (e.g., the claim that 5G telecommunications networks activated the virus), causing a significant negative impact on society Douglas (2021). As such, there is an increasing urgency for high-performance methods that can automatically detect conspiracy theories.
@@
It has previously been shown that there is a close relationship between misinformation (including conspiracy theories) and affective information, i.e., sentiment and emotionsLiu et al. (2024b). For example, Dong et al. (2020) found that during the COVID-19 pandemic, there was a correlation between the level of public anger and the likelihood of rumor propagation, while Zaeem et al. (2020) also observed a significant positive correlation between negative emotions and fake news. Based on such observations, many studies have adopted affective information as the means to detect misinformation Liu et al. (2023); Zhang et al. (2021). Here, we aim to extend the study of affective information, specifically sentiment and emotions, to deepen our understanding and detection of conspiracy theories.
@@
Pre-trained language models (PLMs) such as BERT Devlin et al. (2018) and RoBERTa Liu et al. (2019) have shown outstanding performance when applied to various classification tasks, including conspiracy theory detection Yanagi et al. (2021); Peskine et al. (2021). However, due to their restricted number of parameters. PLMs do not perform optimally when applied to diverse and complex tasks Zhang et al. (2023). Recently, LLMs, which possess significantly larger numbers of parameters, have been explored as a novel means to address the issue of misinformation, with very promising results Hu et al. (2023); Pavlyshenko (2023); Cheung and Lam (2023). However, these studies mostly focus on binary classification of texts according to whether or not they convey misinformation, Moreover, these previous studies have mostly utilized simple prompts to test or carry out instruction-tuning of LLMs, or else have employed LLMs as auxiliary tools for other models. To our knowledge, no existing LLM-based studies have attempted either to leverage important affective features that are characteristic of misinformation, or are able to carry out in-depth analyses of conspiracy-related text.
@@
To address these research gaps, we have constructed a multitask conspiracy detection instruction dataset, ConDID, to facilitate instruction-tuning and evaluation of LLMs. Based on annotations in two conspiracy theory datasets, ConDID is divided into five tasks that encompass conspiracy theory judgment, conspiracy theory topic detection, and conspiracy theory intention detection. We subsequently propose a novel open-source LLM, ConspEmoLLM, that is specialized for detecting conspiracy-related information. ConspEmoLLM is created by applying an instruction-tuning method to an emotion-oriented LLM using the ConDID dataset. Evaluation of ConspEmoLLM using the ConDID test set shows that it achieves state-of-the-art (SOTA) performance among other open-source LLMs, as well as the closed-source ChatGPT.
@@
(2) We conduct an affective analysis on the two conspiracy theory datasets used to construct ConDID, which provides evidence that expressions of sentiment and emotions in conspiracy theory text are distinct from those occurring in mainstream text.
@@
The remainder of this paper is structured as follows: Section 2 introduces related work concerning conspiracy theory and misinformation detection; sentiment analysis; and open-source LLMs. Section 3 describes our construction of the ConDID dataset, the affective analysis of conspiracy theory datasets, and the instruction-tuning of ConspEmoLLM. Section 4 reports our evaluation and analysis of the performance of multiple models on the ConDID test set. Section 5 concludes the paper and provides directions for future work. Section 6 and Section 7, respectively, discuss potential limitations and confirm the ethical soundness of our study. Appendix A provides figures illustrating our affective analysis of conspiracy theory datasets.
@@ 2.1 Conspiracy theory and misinformation detection
PLMs have been widely applied to the task of conspiracy theory and misinformation detection. For example, Yanagi et al. (2021) utilized BERT as the base model for COVID-19 conspiracy theory detection, while Peskine et al. (2021) applied a domain-specific COVID BERT (CT-BERT) to the same task. More recently, an increasing amount of attention has been devoted to exploring the application of LLMs to detect conspiracy theories and misinformation. For example, Peskine et al. (2023) employed zero-shot learning to evaluate the accuracy of the GPT-3 model in performing fine-grained multi-label conspiracy theory classification of tweets, while Hu et al. (2023) proposed a fake news detection framework that utilizes an LLM as an auxiliary tool to enhance the prediction accuracy of BERT. Meanwhile, Pavlyshenko (2023) employed prompt-based fine-tuning of LLaMA2 for rumor and fake news detection. Cheung and Lam (2023) supplemented an LLM with external knowledge to enhance the performance of fake news detection. However, all of these models focus on binary classification and do not exploit affective information for misinformation detection.
@@ 2.2 Affective analysis
Automated affective analysis of text has previously been carried out by various means, including the use of different sentiment analysis tools, such as VADER Hutto and Gilbert (2014) and TextBlob [^1] https://textblob.readthedocs.io/. PLMs have also been employed for sentiment analysis. For example, Hoang et al. (2019) used BERT for aspect-based sentiment analysis, while Tan et al. (2022) proposed a hybrid sentiment analysis model that combines RoBERTa with an LSTM. More recent work has also begun to explore the effectiveness of LLMs for sentiment analysis. For instance, Zhang et al. (2023) and Lei et al. (2023) both utilized retrieval-augmented LLMs to enhance the sentiment analysis capabilities of LLMs when applied to financial news and dialogues, respectively. Meanwhile, Liu et al. (2024a) proposed a series of comprehensive LLMs that are specialized for affective analysis (EmoLLMs), and which are capable of analyzing emotions across five different dimensions (i.e. emotion intensity, ordinal classification of emotion intensity, sentiment strength, sentiment classification, emotion detection). EmoLLMs demonstrate strong generalization ability, surpassing ChatGPT and GPT-4 in most emotion analysis tasks. Therefore, we use one of these EmoLLMs, i.e., EmoLLaMA, to perform affective analysis in this study.
@@ 2.3 Open-source LLMs
A large amount of research has been dedicated to developing open-source LLMs as an alternative to the well-known closed-source LLMs (e.g., ChatGPT), in order to support easier research into the improvement and application of LLMs. Popular series of open-source, general language models include LLaMA Touvron et al. (2023), OPT Zhang et al. (2022) and BLOOM Workshop et al. (2022). These are complemented by a range of domain-specific open-source LLMs, including FinMA Xie et al. (2023) for finance, MentalLLaMA Yang et al. (2023) for mental health, ExTES-LLaMA Zheng et al. (2023) for emotional support chatbots, and TimeLLaMA Yuan et al. (2023) for temporal reasoning. In this work, we extend the inventory of domain-specific LLMs, by developing the first open-source LLM for multitask conspiracy theory detection based on affective information.
@@ 3.2.1 Raw data
COCO The COVID-19 conspiracy theories (COCO) dataset Langguth et al. (2023) is an extension of the dataset used in the MediaEval FakeNews: Corona Virus and Conspiracies Task challenge Pogorelov et al. (2021). COCO consists of tweets, each of which is assigned 12 different labels that characterize the intention of the tweet with respect to 12 different conspiracy theory categories  [^2] Suppressed Cures, Behavior Control, Anti Vaccination, Fake Virus, Intentional Pandemic, Harmful Radiation, Depopulation, New World Order, Esoteric Misinformation, Satanism, Other Conspiracy Theory, Other Misinformation.. Each label can have three possible values, i.e., Unrelated: the tweet is unrelated to the specific conspiracy category. This type of tweet contains keywords related to the conspiracy, but they are used in a completely different context; Related: The tweet is related to the specific category, but it does not propagate misinformation or conspiracy theories; Conspiracy: The tweet is related to the specific category and is actively aimed at spreading conspiracy theories.
Each tweet in COCO is also assigned a single overall intention label, as follows: Conspiracy is assigned to tweets for which the Conspiracy label is assigned to at least one of the 12 categories. Otherwise, if the Related category is assigned to at least one of the categories, then the overall label of Related is used. The overall label of Unrelated is only used for tweets that are unrelated to all 12 conspiracy categories.
LOCOAnnotations The Language of Conspiracy (LOCO) Miani et al. (2021) is corpus consisting of documents gathered from the internet (88 million words), which is used to study the differences between conspiracy language and mainstream language. We use an annotated subset of LOCO, created by Mompelat et al. (2022), which we refer to as LOCOAnnotations. This subset consists of documents concerning two different topics (i.e., the Sandy Hook school shooting and coronavirus) in which two types of labels have been assigned. The first type of label concerns whether or not the document is directly concerned with a conspiracy theory. The second type of label reflects the degree of relatedness to a conspiracy theory, using three labels (closely related/broadly related/not related).
@@ 3.2.2 Tasks
Using these two different datasets, we define five different tasks. Tasks 1-3 are based on the COCO corpus, and are similar to those used in the MediaEval FakeNews challenge, while Tasks 4 and 5 and based on LOCOAnnotations.
@@
Task 2: Conspiracy Theory Topics Detection Determine whether the tweet mentions or refers to any predefined conspiracy theory categories.
Task 3: Combination of Task 1 and Task 2 Predict both the conspiracy category and the relationship of the tweet (Unrelated/Related/Conspiracy) with the category.
@@ 3.2.3 Affective analysis of raw data
EmoLLMs Liu et al. (2024a) is a series of affective analysis models. We adopt the best-performing model from this series (i.e., EmoLLaMA-chat-7b) to conduct sentiment analysis on COCO and LOCOAnnotations across five different dimensions (i.e., emotion intensity, ordinal classification of emotion intensity, sentiment strength, sentiment classification, emotion detection). The results of this analysis are shown in Figures 1 to 5. Due to space limitations, these figures only concern the analysis of intentions in the COCO dataset; sentiment analysis figures for different conspiracy categories in COCO dataset and LOCOAnnoations, can be found in the Appendix A.1 and Appendix A.2. Figures 1 to 5 depict the differences between affective features according to the intentions of tweets (i.e., Unrelated/Related/Conspiracy). Meanwhile, Figures 7 to 11 in Appendix A.1 illustrate the affective differences according to the 12 different conspiracy categories. Figures 12 to 21 in Appendix A.2 compare affective differences between conspiracy and non-conspiracy text, and text exhibiting various levels of relatedness to conspiracy theories (i.e., Not related/closely related/broadly related).
@@
Figure 2: Ordinal classification of emotion intensity of different intentions
@@
It can be observed from Figures 1 to 5 that tweets related to conspiracy theories predominantly convey negative sentiments and emotions (e.g., anger, fear, and disgust). In contrast, the Unrelated tweets are more likely to express positive sentiments and emotions (e.g., joy, love, and optimism). Figures 7 to 11 reveal that the types of emotions expressed in tweets can also vary according to the specific category of conspiracy theory being discussed. In the LOCOAnnotations dataset, there are also noticeable differences in the types of emotion information conveyed across different categories and levels of conspiracy relatedness.

### v3

@@ Abstract
The internet has brought both benefits and harms to society. A prime example of the latter is misinformation, including conspiracy theories, which flood the web. Recent advances in natural language processing, particularly the emergence of large language models (LLMs), have improved the prospects of accurate misinformation detection. However, most LLM-based approaches to conspiracy theory detection focus only on binary classification and fail to account for the important relationship between misinformation and affective features (i.e., sentiment and emotions). Driven by a comprehensive analysis of conspiracy text that reveals its distinctive affective features, we propose ConspEmoLLM, the first open-source LLM that integrates affective information and is able to perform diverse tasks relating to conspiracy theories. These tasks include not only conspiracy theory detection, but also classification of theory type and detection of related discussion (e.g., opinions towards theories). ConspEmoLLM is fine-tuned based on an emotion-oriented LLM using our novel ConDID dataset, which includes five tasks to support LLM instruction tuning and evaluation. We demonstrate that when applied to these tasks, ConspEmoLLM largely outperforms several open-source general domain LLMs and ChatGPT, as well as an LLM that has been fine-tuned using ConDID, but which does not use affective features. ConspEmoLLM can be easily applied to identify and classify conspiracy-related text in the real world. The work has been released at https://github.com/lzw108/ConspEmoLLM/.
[…]
123
@@ 1 Introduction
Misinformation has become one of the major threats to society. The rise of the internet and social media has made it increasingly simple for misinformation to spread rapidly. Conspiracy theories are one type of misinformation, whose false content is intended to cause harm [19]. Examples of popular conspiracy theories include those claiming that the Earth is flat and that vaccines cause autism. Conspiracy theorists ignore scientific evidence and tend to interpret events as secretive actions [5]. During the COVID-19 pandemic, the spread of conspiracy theories significantly increased (e.g., the claim that 5G telecommunications networks activated the virus), resulting in a considerable negative impact on society [4]. As such, there is an increasing urgency for high-performance methods that can automatically detect conspiracy theories.
@@
It has previously been shown that there is a close relationship between misinformation (including conspiracy theories) and affective information, i.e., sentiment and emotions [14]. For example, Dong et al. [3] found that during the COVID-19 pandemic, there was a correlation between the level of public anger and the likelihood of rumor propagation, while Zaeem et al. [33] also observed a significant positive correlation between negative emotions and fake news. Based on such observations, many studies have adopted affective information as the means to detect misinformation [11, 36]. Here, we aim to extend the study of affective information, specifically sentiment and emotions, to deepen our understanding of conspiracy theories and to improve their automated detection.
@@
Pre-trained language models (PLMs) such as BERT [2] and RoBERTa [12] have shown outstanding performance when applied to various classification tasks, including conspiracy theory detection [30, 22]. However, due to their restricted number of parameters, PLMs cannot perform optimally when applied to diverse and complex tasks [34]. Recently, LLMs, which possess significantly larger numbers of parameters, have been explored as a novel means to address the issue of misinformation, with very promising results [7, 20, 1]. However, these studies mostly focus on binary classification of texts according to whether or not they convey misinformation, Moreover, these previous efforts have mostly utilized simple prompts to test or carry out instruction-tuning of LLMs, or else have employed LLMs as auxiliary tools for other models. To our knowledge, no existing LLM-based studies have attempted to leverage important affective features that are characteristic of misinformation, nor have such studies attempted to carry out in-depth analyses of conspiracy-related text.
@@
To address these research gaps, we have constructed a multitask conspiracy detection instruction dataset, ConDID, to facilitate instruction-tuning and evaluation of LLMs. Based on annotations in two conspiracy theory datasets, ConDID is divided into five tasks that encompass conspiracy theory judgment, conspiracy theory topic detection, and conspiracy theory intention detection. We subsequently propose a novel open-source LLM, ConspEmoLLM, which is specialized for detecting conspiracy-related information. ConspEmoLLM is created by applying an instruction-tuning method to an emotion-oriented LLM using the ConDID dataset. Evaluation of ConspEmoLLM using the ConDID test set shows that it achieves state-of-the-art (SOTA) performance among other open-source LLMs, as well as the closed-source ChatGPT.
@@
(2) We conduct an affective analysis of the two conspiracy theory datasets used to construct ConDID, which provides evidence that expressions of sentiment and emotions in conspiracy theory text are distinct from those occurring in the mainstream text.
@@
The remainder of this paper is structured as follows: Section 2 introduces related work concerning the detection of conspiracy theories and misinformation, analysis of sentiment, and open-source LLMs. Section 3 describes our construction of the ConDID dataset, the affective analysis of conspiracy theory datasets, and the instruction-tuning of ConspEmoLLM. Section 4 reports on our evaluation and analyzes of the performance of multiple models on the ConDID test set. Section 6 concludes the paper and provides directions for future work. Sections 7 and 8, respectively, discuss potential limitations and confirm the ethical soundness of our study. The supplementary material provides figures illustrating our detailed affective analysis of conspiracy theory datasets.
@@ 2.1 Conspiracy theory and misinformation detection
PLMs have been widely applied to the task of conspiracy theory and misinformation detection. For example, Yanagi et al. [30] utilized BERT as the base model for COVID-19 conspiracy theory detection, while Peskine et al. [22] applied a domain-specific COVID BERT (CT-BERT) to the same task. More recently, an increasing amount of attention has been devoted to exploring the application of LLMs to detect conspiracy theories and misinformation. For example, Peskine et al. [23] employed zero-shot learning to evaluate the accuracy of the GPT-3 model in performing fine-grained multi-label conspiracy theory classification of tweets, while Hu et al. [7] proposed a fake news detection framework that utilizes an LLM as an auxiliary tool to enhance the prediction accuracy of BERT. Meanwhile, Pavlyshenko [20] employed prompt-based fine-tuning of LLaMA2 for rumor and fake news detection. Cheung and Lam [1] supplemented an LLM with external knowledge to enhance the performance of fake news detection. However, all of these models focus on binary classification and do not exploit affective information for misinformation detection.
@@ 2.2 Affective analysis
Emotion detection and sentiment analysis are two types of NLP techniques for analyzing human expressions. Sentiment analysis aims to capture both the overall emotional tone conveyed by the data source (typically positive, negative, or neutral) and the intensity of this tone. Emotion detection is the process of categorizing data at a finer level of granularity based on the emotions conveyed. In comparison to sentiments, emotions correspond to more specific and intense feelings. For example, negative emotions include anger and fear, while positive emotions include happiness and joy [14]. Identifying both sentiments and emotions is crucial for downstream tasks.
Automated affective analysis of text has previously been carried out by various means, including the use of different sentiment analysis tools, such as VADER [8] and TextBlob [^1] https://textblob.readthedocs.io/. PLMs have also been employed for sentiment analysis. For example, Hoang et al. [6] used BERT for aspect-based sentiment analysis, while Tan et al. [26] proposed a hybrid sentiment analysis model that combines RoBERTa with an LSTM. More recent work has also begun to explore the effectiveness of LLMs for sentiment analysis. For instance, Zhang et al. [34] and Lei et al. [10] both utilized retrieval-augmented LLMs to enhance the sentiment analysis capabilities of LLMs when applied to financial news and dialogues, respectively. Meanwhile, Liu et al. [13] proposed a series of comprehensive LLMs that are specialized for affective analysis (EmoLLMs), and which are capable of analyzing emotions across five different dimensions (i.e. emotion intensity, ordinal classification of emotion intensity, sentiment strength, sentiment classification, emotion detection). EmoLLMs demonstrate strong generalization ability, surpassing ChatGPT and GPT-4 in most emotion analysis tasks. Therefore, we use one of these EmoLLMs, i.e., EmoLLaMA, to perform affective analysis in this study.
@@ 2.3 Open-source LLMs
A large amount of research has been dedicated to developing open-source LLMs as an alternative to the well-known closed-source LLMs (e.g., ChatGPT), in order to support easier research into the improvement and application of LLMs. Popular series of open-source, general language models include LLaMA [27], OPT [35] and BLOOM [28]. These are complemented by a range of domain-specific open-source LLMs, including FinMA [29] for finance, MentalLLaMA [31] for mental health, ExTES-LLaMA [37] for emotional support chatbots, and TimeLLaMA [32] for temporal reasoning. In this work, we extend the inventory of domain-specific LLMs, by developing the first open-source LLM for multitask conspiracy theory detection based on affective information.
@@ 3.2.1 Raw data
COCO The COVID-19 conspiracy theories (COCO) dataset [9] is an extension of the dataset used in the MediaEval FakeNews: Corona Virus and Conspiracies Task challenge [24]. COCO consists of tweets, each of which is assigned 12 different labels that characterize the intention of the tweet with respect to 12 different conspiracy theory categories  [^2] Suppressed Cures, Behavior Control, Anti Vaccination, Fake Virus, Intentional Pandemic, Harmful Radiation, Depopulation, New World Order, Esoteric Misinformation, Satanism, Other Conspiracy Theory, Other Misinformation.. Each label can have three possible values, i.e., Unrelated: The tweet is unrelated to the specific conspiracy category; it contains conspiracy-related keywords, but they are used in a completely different context; Related: The tweet is related to the specific category, but it does not propagate misinformation or conspiracy theories; Conspiracy: The tweet is related to the specific category and is actively aimed at spreading conspiracy theories.
Each tweet in COCO is also assigned a single overall intention label, as follows: The overall Conspiracy label is assigned to tweets for which the Conspiracy label is used for at least one of the 12 categories. Otherwise, if the Related category is assigned to at least one of the categories, then the overall label of Related is used. The overall label of Unrelated is only used for tweets that are unrelated to all 12 conspiracy categories.
LOCOAnnotations The Language of Conspiracy (LOCO) [16] is a corpus consisting of documents gathered from the internet (88 million words), which is used to study the differences between conspiracy language and mainstream language. We use an annotated subset of LOCO, created by Mompelat et al. [17], which we refer to as LOCOAnnotations. This subset consists of documents concerning two different topics (i.e., the Sandy Hook school shooting and coronavirus) in which two types of labels have been assigned. The first type of label concerns whether or not the document is directly concerned with a conspiracy theory. The second type of label reflects the degree of relatedness to a conspiracy theory, using three labels (closely related, broadly related or not related).
@@ 3.2.2 Tasks
Using these two different datasets, we define five different tasks. Tasks 1-3 are based on the COCO corpus, and are similar to those used in the MediaEval FakeNews challenge, while Tasks 4 and 5 are based on LOCOAnnotations.
@@
Task 2: Conspiracy Theory Topics Detection Determine whether the tweet mentions or refers to any of the 12 predefined conspiracy theory categories.
Task 3: Combination of Task 1 and Task 2 Predict both the conspiracy category and the relationship of the tweet (Unrelated/Related/Conspiracy) to the category.
@@ 3.2.3 Affective analysis of raw data
EmoLLMs [13] is a series of affective analysis models. We adopt the best-performing model from this series (i.e., EmoLLaMA-chat-7b) to conduct sentiment analysis on COCO and LOCOAnnotations across five different dimensions, as follows:
  • Emotion intensity score: For each of four different emotions (anger, fear, joy and sadness), a real-valued score between 0 and 1 is assigned to represent the intensity of the emotion in the text.
  • Emotion intensity classification: For the same four emotions, one of four different emotional intensity classes is assigned to the text, i.e. {no, low, moderate, high} emotional intensity.
  • Sentiment strength score: A real-valued score between 0 (most negative) and 1 (most positive) is assigned to represent the intensity of sentiment expressed in the text.
  • Sentiment classification: One of seven classes (i.e. {very, moderately, slightly} negative, neutral, {slightly, moderately, very} positive) is assigned to represent the intensity of sentiment conveyed in the text.
  • Emotion detection: One or more labels is assigned to the text if any of eleven different emotions are conveyed (i.e. anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise and trust). If none of these emotions is expressed, then the label neutral or no emotion is assigned.
The results of analyzing the COCO dataset according to these five dimensions are shown in Figures 1 to 5, broken down according to the different intentions of the tweets in the dataset (i.e., Unrelated/Related/Conspiracy). In Figures 1 and 3, the x-axis represents the real-valued scores for intensity of different emotions and sentiment strength, respectively. The y-axis represents the corresponding probability density distribution (i.e., the number of tweets belonging to each intention class that has a particular score for emotional intensity/sentiment strength). In Figures 2, 4 and 5, the y-axis represents the distribution of labels within the intention class indicated on the x-axis. Due to space limitations, these figures only report on part of our complete analysis. The supplementary material includes additional figures that depict an analysis of the five dimensions based on the different conspiracy categories of tweets in the COCO dataset. The supplementary material also includes analyses of LOCOAnnoations, which examine affective differences between conspiracy and non-conspiracy text, and between text exhibiting various levels of relatedness to conspiracy theories (i.e., Not related/closely related/broadly related).
@@
Figure 2: Emotion intensity classification of different intentions
@@
Figure 1 presents a quantitative analysis of emotional intensity. It may be observed that tweets that refer to conspiracy theories (i.e., Related/Conspiracy) express stronger feelings of anger and fear, with a lesser degree of joy, compared to the Unrelated tweets. The intensity of sadness is similar across the three categories, indicating that during the pandemic, people were generally in a state of sadness. Similar conclusions may be drawn from Figure 2, which presents a qualitative analysis of emotional intensity. The sentiment strength in Figure 3 and the sentiment polarity classification in Figure 4 indicate that sentiments expressed in Unrelated tweets are more strongly positive than sentiments in tweets related to conspiracy theories. It can be observed from Figure 5 that tweets related to conspiracy theories predominantly convey negative sentiments and emotions (e.g., anger, fear, and disgust). In contrast, the Unrelated tweets are more likely to express positive sentiments and emotions (e.g., joy, love, and optimism). The figures in the supplementary materials reveal that the types of emotions expressed in tweets can also vary according to the specific category of conspiracy theory being discussed in the COCO dataset, and that there are noticeable differences in the types of emotion information conveyed across different categories and levels of conspiracy relatedness in the LOCOAnnotations dataset. Overall, it can be inferred that there are close relationships between affective information and conspiracy theories. The affective information can be an important feature for detecting conspiracy theories.